OpenStudy (anonymous):

\[\int\limits_{-1}^{0}{2 \over 3x-5}dx\] I think I know how to integrate this, but my substitution isn't working?

OpenStudy (anonymous):

Remember the general form, on the picture

OpenStudy (anonymous):

Yes, isn't the integral form 2/3 * ln(3x-5)?

OpenStudy (anonymous):

yes

sam (.sam.):

int[-1,0,(2)/(3x-5),x] let u=3x-5 dx=(du)/(3) (2ln(|u|))/(3) (2ln(|(3(0)-5)|))/(3)-[(2ln(|(3(-1)-5)|))/(3)] ln((5^(2)/(3)*(1)/(4)

OpenStudy (anonymous):

But how do you substitute the 0 and -1?

sam (.sam.):

Integrate with respect to u, then substitute back u=3x-5

sam (.sam.):

Then only use 0 and -1

OpenStudy (anonymous):

since you have your integration, you will want to plug in the upper bound x value which is 0 to the integration. And whatever you get subtract from it the value you get by pluging in the lower bound -1 to the integration.

OpenStudy (anonymous):

Yes, I know that, but I end up with 2/3 * In(-5) - In(-8) and I don't know how to carry on from there..?

OpenStudy (anonymous):

If it comes out to that then probably there is no solution

OpenStudy (anonymous):

but the answer is 2/3In5 - In4 ~~~in my textbook

OpenStudy (anonymous):

are you sure it is ln(3x - 5)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thats wierd

OpenStudy (anonymous):

i will redo it and see

sam (.sam.):

\[\huge \int\limits_{-1}^{0}\frac{2}{3x-5}dx\] Let u=3x-5 du=3dx \[\frac{du}{3}=dx\] Substitute du/3 to dx \[\huge \int\limits\limits_{-1}^{0}\frac{2}{u}\frac{du}{3}\] \[\huge \frac{2}{3}\int\limits\limits\limits_{-1}^{0}\frac{1}{u}du\] \[\huge \frac{2}{3}[\ln(u)]^{0}_{-1}\] u=3x-5 \[\huge \frac{2}{3}[\ln(3x-5)]^{0}_{-1}\] \[\huge \frac{2}{3}[\ln(-5)-\ln(-8)]\]

OpenStudy (anonymous):

I am still getting the same result, maybe that was a typo

OpenStudy (anonymous):

Thanls Sam. That's what I got... but the answer in the textbook was different. Hmmm. Perhaps..?

OpenStudy (anonymous):

Or the textbook is wrong

OpenStudy (anonymous):

No typo. I know. I have the textbook right infront of me.

OpenStudy (anonymous):

Are you learning on your own or is this a class

sam (.sam.):

Wait for the natural log, inside is absolute

OpenStudy (anonymous):

I'm learning on my own :(

OpenStudy (anonymous):

That is amazing, keep it up. :) If you dont mind how old are you

OpenStudy (anonymous):

I can't say my age... but I'm trying to do 2 years of A-level in one for 4 subjects... And the school doesn't combine 2 years teaching in one, so I have to do 1 of the two years on my own...

sam (.sam.):

\[\huge \int\limits_{-1}^{0}\frac{2}{3x-5}dx\] Let u=3x-5 du=3dx \[\frac{du}{3}=dx\] Substitute du/3 to dx \[\huge \int\limits\limits_{-1}^{0}\frac{2}{u}\frac{du}{3}\] \[\huge \frac{2}{3}\int\limits\limits\limits_{-1}^{0}\frac{1}{u}du\] \[\huge \frac{2}{3}[\ln(u)]^{0}_{-1}\] u=3x-5 \[\huge \frac{2}{3}[\ln|3x-5|]^{0}_{-1}\] \[\huge \frac{2}{3}[\ln5-\ln8]\] \[\huge \frac{2}{3}[\ln\frac{5}{8}]\]

OpenStudy (anonymous):

Sam, thanks so much for trying, but that doesn't seem to work as the answer either :/

OpenStudy (anonymous):

maybe the book made a mistake?

sam (.sam.):

whats the answer

OpenStudy (anonymous):

Keep it up, i will also suggest that you use www.khanacademy.org he has free calculus tutorials.

OpenStudy (anonymous):

2/3In5 - In4

OpenStudy (anonymous):

Thanks deny!

sam (.sam.):

\[\huge \frac{2}{3}[\ln5-\ln8]\] \[\huge \frac{2}{3}\ln5-\frac{2}{3}\ln8\] \[\frac{2}{3}\ln5-\ln8^{\frac{2}{3}}\] \[\frac{2}{3}\ln5-\ln4\]

OpenStudy (anonymous):

Ok, that's the only explanation :) Thanks!

sam (.sam.):

no problem :)

OpenStudy (anonymous):

@.Sam. good job, I didn't even think about that